Let {an}∞n=1 be a convergent sequence with limit L. Then prove that any subsequence of {an}∞n=1 is also convergent and has limit L


1
Expert's answer
2021-02-19T15:12:26-0500

Given that "\lim \limits_{n \to \infty} a_n = L" . This implies that given "\epsilon >0, \exist N \in \mathbb{N} \ni | a_n - L| < \epsilon" whenever "n \geq N"

Let "a_{n_k}" be a subsequence of "a_n" . Since "n_k \geq n" , we have that "n_k \geq N"

Consider,

"|a_{n_k} - L| = | a_{n_k} - a_n + a_n - L|\\ \leq |a_{n_k} - a_n| + | a_n - L|"

Since every convergent sequence is a Cauchy sequence we have that "|a_{n_k} - a_n| < \epsilon"

Also, since "a_n" is convergent we have that "|a_n - L| < \epsilon"

So, we have that "|a_{n_k} - L | < \epsilon + \epsilon = 2\epsilon"

By the "K-\epsilon" Lemma, we have that "a_{n_k}" is convergent, and converges to L.